Equivalent weight of carbon monoxide and carbon dioxide in the ratio of

Grade:12th pass

1 Answers

Arun
6 years ago
 
Dear student
Equivalent weight of C in CO2 is 
12g/4= 3 g
Because the valency of C is 4 here.
 
Similarly equivalent weight of C in CO is
12g/2 = 6 gm
Because the valency of C is 2 here.
 
Hence ratio of equivalent weight of C in CO2 & CO=
= 3/6 = 1/2
= 1 :2
 
Regards
Arun (askIITians forum expert)